OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x + y – 3z = –8 2x + 2y + z = 12 3x + y – z = –2 Choices: A.x + y = –8 3x + y = –2 B.7x + 7y = 28 5x + 3y = 10 C.x + y = –8 5x + 3y = 14 D.7x + 7y = 28 3x + y = –2

OpenStudy (sheetal):

B

OpenStudy (sheetal):

x+y-3z=-8)*1 2x+2y+z=12)*-3 (jst to make coefficient of z equal) =x+y-3z=-8 -6x-6y-3z=-36 + + + + =7x+7y= 28 similarly 2x+2y+z=12)*-1 3x+y-z=-2)*1 = -2x-2y-z=-12 3x+y-z=-2 - - + = -5x-3y=-10 = -(5x+3y)=-10 =5x+3y=10